# Sequential Monte-Carlo testing by betting

### v4

@@ 1.3 Outline of the paper and our contributions
In this paper we construct anytime-valid permutation tests based on novel test martingales. Our proposed e-values and p-values have easy calculable closed forms, save computational efforts under both the null and the alternative and can bound the resampling risk at arbitrary small values. We start with two simple betting strategies which serve to exemplify the approach and provide some first intuition of the opportunities it offers (Section 2). Afterwards, we derive the log-optimal betting strategy and show that the obtained wealth does not grow exponentially (Section 3). Based on these results, we introduce in Section 4 a simple strategy that is based on a single parameter indicating aggressiveness and which leads to a wealth that is proportional to the likelihood of a binomial distribution. In Section 5, we show how the log-optimal strategy for a given prior can be obtained as a mixture of these simple binomial strategies and use this to introduce a certain prior under which the strategy always rejects when the limiting permutation p-value $\mathfrak{p}^{\mathrm{lim}}$ is less than $c$ for some constant $c<\alpha$ arbitrary close to $\alpha$. In Section 6, we show that every rank-based permutation test can be obtained by our general betting approach. In particular, we derive concrete betting strategies that yield anytime-valid generalizations of the classical permutation p-value and the Besag-Clifford method. We compare our approach to the Besag-Clifford strategy [3] and the classical permutation p-value via simulations and real data analyses in Section 7. The R code to reproduce all figures and results is available at the GitHub repository github.com/fischer23/MC-testing-by-betting.
@@ 2 The aggressive strategy
The most passive betting strategy uses the betting vector
$1,1,\dots,1$ (3)
meaning that $B_{t}(r)=1$ for all $r$ and thus no matter the outcome, our wealth is multiplied by 1. Thus, the wealth starts at one and stays at 1, meaning that the resulting test is powerless.
In contrast, the most aggressive betting strategy uses the vector
$0,\frac{t+1}{t},\frac{t+1}{t},\dots,\frac{t+1}{t},$ (4)
meaning that it sets $B_{t}(1)=0$ and $B_{t}(r)=(t+1)/t$ for all other $r$. This would be the strategy of a gambler who not only strongly believes that $H_{1}$ is true, but also believes that $Y_{0}$ is the largest amongst all other $Y_{i}$’s, and so the gambler deems it is (nearly) impossible for $Y_{t}$ to have the smallest rank, forsaking all his money if that happens. Since our alternative $H_{1}$ states that $Y_{t}$ is exchangeable with $Y_{1},\dots,Y_{t-1}$ conditional on $Y_{0}$, its rank among those is going to be uniform, and thus it makes sense to have $B_{t}(r)=B_{t}(r^{\prime})$ for any $r\neq 1,r^{\prime}\neq 1$ (We underpin this statement theoretically in Section 3). This is a risky strategy: the first round $t$ in which $Y_{t}$ ends up larger than $Y_{0}$, we would multiply our wealth by 0 and can never recover from it. However, if our gamble was actually always correct, then our wealth after $B$ rounds would be $\prod_{t=1}^{B}(t+1)/t=B+1$, and this is the largest possible wealth. Indeed, since the inverse of the wealth is a valid anytime-valid p-value, it would yield a p-value of $1/(B+1)$, the best possible in this case.
The aggressive gambling strategy may initially not seem too prudent. It may appear “too aggressive”, and in many real-world problems will end up with zero wealth, a seemingly disappointing outcome. While this is true, the final wealth does not necessarily convey the full story. Recall that
$\mathfrak{p}^{\mathrm{agg}}_{t}:=(\max_{s\leq t}W_{s})^{-1}$
defines a p-process, yielding a p-value at any stopping time. Let us now examine this p-value at certain natural stopping times. Define the stopping time
$\tau:=\inf\{t\geq 1:W_{t}=0\}=\inf\{t\geq 1:Y_{t}\geq Y_{0}\},$ (5)
which is almost surely finite [^1] It may appear as if $\tau$ could hypothetically be infinite if the strategy never loses and it is never stopped, but if one is recalculating the test statistic on uniformly randomly drawn permutations (as is the norm in practice), $\tau$ will be almost surely finite because the identity permutation will certainly be drawn at some point with probability one. In fact, the distribution of $\tau$ is a (shifted) geometric distribution, with finite mean and variance.. Interestingly, a short calculation reveals that
$\mathfrak{p}^{\mathrm{agg}}_{\tau}=1/\tau.$
@@ 2.1 Connection to Besag-Clifford
The aggressive strategy yields an interesting connection to the sequential permutation test by Besag and Clifford [3], when their method is stopped after one loss, where a “loss” refers to the event $\{Y_{t}\geq Y_{0}\}$ for some $t>0$. That means the (random) number of losses after $B$ permutations equals $K_{B}:=\mathrm{rank}(Y_{0};Y_{0}^{B})-1$. Define the stopping time $\gamma(h,B)$ that stops after $h$ losses or after sampling $B$ permutations, whichever happens first, and let
$\gamma(h):=\gamma(h,\infty)$
denote the stopping time when we do not impose any finite maximum stopping time. Besag and Clifford considered the case $B<\infty$ and defined a single p-value $\mathfrak{p}^{\mathrm{BC}}_{\gamma(h,B)}$ (as opposed to a p-process) as
$\mathfrak{p}^{\mathrm{BC}}_{\gamma(h,B)}=\begin{cases}h/\gamma(h,B),&K_{\gamma(h,B)}=h\\ (K_{\gamma(h,B)}+1)/(B+1),&\text{otherwise}.\end{cases}$
@@
The p-value $\mathfrak{p}^{\mathrm{BC}}_{\gamma(h,B)}$ is valid only at time $\gamma(h,B)$ and not at any other stopping time. Since our test yields a p-value at any stopping time, it makes sense to ask what it equals at time $\gamma(1,B)$. The answer is that our p-process equals the Besag-Clifford p-value at that stopping time:
$\mathfrak{p}^{\mathrm{agg}}_{\gamma(1,B)}=\mathfrak{p}^{\mathrm{BC}}_{\gamma(1,B)}.$
@@
The above equality has two key implications. First, it clearly leads to the interpretation that our aggressive betting strategy yields a p-process that is an anytime-valid generalization of the Besag-Clifford p-value that stops after one loss. (Our p-process recovers the latter p-value exactly at time $\gamma(1,B)$, but remains valid at all other stopping times.) Second, it is well known that $\mathfrak{p}^{\mathrm{BC}}_{\gamma(1,B)}$ is in fact an exact p-value, meaning that even though it can only take on the values $1/1,1/2,1/3,\dots,1/(B+1)$, we have $\mathbb{P}(\mathfrak{p}^{\mathrm{BC}}_{\gamma(1,B)}\leq c)=c$ for any $c$ in this support. Thus, the same property also holds for $\mathfrak{p}^{\mathrm{agg}}_{\gamma(1,B)}$.
2.2 The case of $B=\infty$: a negative binomial permutation test
It turns out that the Besag-Clifford strategy also offers an exact p-value if we do not specify an upper bound $B$ of maximum permutations, meaning that the p-value
$\mathfrak{p}_{\gamma(h)}:=\lim_{B\to\infty}\mathfrak{p}^{\mathrm{BC}}_{\gamma(h,B)}$
is also valid and exact. While we do not know of a reference for the case $B=\infty$, a direct argument of this fact follows the same argument as the $B<\infty$ case by Besag and Clifford, which we provide below as a self-contained result for easier later reference. The maximum number of permutations drawn is still almost surely finite, and has finite expectation and variance. Indeed, when the permutations are sampled with replacement, $\gamma(r)$ is just a sum of (shifted) geometric random variables, or equivalently, $\gamma(r)-r$ is negative binomial, so we call the test as the negative binomial permutation test.
Proposition 2.1 (Negative binomial permutation test).
For any pre-defined number of losses $h\geq 1$, define $\gamma(h):=\inf\{t\geq 1:\sum_{i=1}^{t}1\{Y_{t}\geq Y_{0}\}\geq h\}$ as being the time of the $h$-th loss. Then, $\mathfrak{p}_{\gamma(h)}:=h/\gamma(h)$ is a valid and exact p-value for $H_{0}$: for any $N\geq 0$, $\mathbb{P}_{H_{0}}(\mathfrak{p}_{\gamma(h)}\leq h/(h+N))=h/(h+N)$.
@@ 2.2 The case of $B=\infty$: a negative binomial permutation test
Note that $\mathrm{supp}(\mathfrak{p}_{\gamma(h)})=\{1,h/(h+1),h/(h+2),\ldots\}$ and for every $N\in\mathbb{N}_{0}$, we have $\mathbb{P}_{H_{0}}(\mathfrak{p}_{\gamma(h)}\leq h/(h+N))=\mathbb{P}_{H_{0}}(\gamma(h)>h+N-1)=h/(h+N)$, where the last equation follows from the fact that $\gamma(h)>h+N-1$ means that $Y_{0}$ is among the $h$ largest test statistics of $Y_{0},Y_{1},\ldots,Y_{h+N-1}$. ∎
We remark that $\mathfrak{p}_{\gamma(h)}$ bears an interesting resemblance to the standard permutation p-value $\mathfrak{p}^{\mathrm{perm}}_{B}$. Both of them equal the number of losses divided by the number of permutations; $\mathfrak{p}^{\mathrm{perm}}_{B}$ fixes the denominator to equal $B$ and lets the numerator be determined by the data, while $\mathfrak{p}_{\gamma(h)}$ fixes the numerator as $h$, and lets the denominator be determined by the data. The fact most relevant for us is that since
$\mathfrak{p}^{\mathrm{agg}}_{\tau}=\mathfrak{p}^{\mathrm{agg}}_{\gamma(1)},$
the same exactness of p-value holds for $\mathfrak{p}^{\mathrm{agg}}_{\tau}$. Said differently, one can obtain a valid and exact p-value as the inverse of the first time at which a permuted statistic exceeds the original unpermuted statistic. We summarize this in another proposition for ease of future reference.

### v5

@@ 1.3 Outline of the paper and our contributions
In this paper we construct anytime-valid permutation tests based on novel test martingales. Our proposed e-values and p-values have easy calculable closed forms, save computational efforts under both the null and the alternative and can bound the resampling risk at arbitrary small values. We start with two simple betting strategies which serve to exemplify the approach and provide some first intuition of the opportunities it offers (Section 2). Afterwards, we derive the (oracle) log-optimal betting strategy, which depends on an unknown parameter (Section 3). Based on these results, we introduce in Section 4 a simple strategy that is based on a single parameter indicating aggressiveness and which leads to a wealth that is proportional to the likelihood of a binomial distribution. In Section 5, we show how the log-optimal strategy can be mimicked by a mixture of these simple binomial strategies and use this to introduce a certain working prior under which the strategy always rejects when the limiting permutation p-value $Q_{\mathrm{lim}}$ is less than $c$ for some constant $c<\alpha$ arbitrary close to $\alpha$. In Section 6, we show that every permutation test based on the number of losses can be obtained by our general betting approach. In particular, we derive concrete betting strategies that yield anytime-valid generalizations of the classical permutation p-value and the Besag-Clifford method. We compare our approach to the Besag-Clifford strategy [3] and the classical permutation p-value via simulations and real data analyses in Section 7. The R code to reproduce all figures and results is available at the GitHub repository github.com/fischer23/MC-testing-by-betting.
@@ 2 The aggressive strategy
We begin by discussing two simple strategies to provide some intuition about our approach. In Sections 35 we then introduce our recommended strategies based on maximizing the expected logarithm of the wealth. The most passive betting strategy uses the betting function
$B_{t}(0)=B_{t}(1)=1\quad(t\in\mathbb{N})$ (6)
and thus no matter the outcome, our wealth is multiplied by 1. Thus, the wealth starts at one and stays at 1, meaning that the resulting test is powerless.
In contrast, the most aggressive betting strategy uses the function
$B_{t}(0)=\frac{t+1}{t}\quad\text{and}\quad B_{t}(1)=0\quad(t\in\mathbb{N}),$ (7)
as long as $L_{t-1}=0$. This would be the strategy of a gambler who not only strongly believes that $H_{1}$ is true, but also believes that $Y_{0}$ is the largest amongst all other $Y_{i}$’s, and so the gambler deems it is (nearly) impossible for $Y_{t}$ to be smaller than $Y_{0}$, forsaking all his money if that happens. This is a risky strategy: the first round $t$ in which $Y_{t}$ ends up larger than $Y_{0}$, we would multiply our wealth by 0 and can never recover from it. However, if our gamble was actually always correct, then our wealth after $T$ rounds would be $\prod_{t=1}^{T}(t+1)/t=T+1$, and this is the largest possible wealth. Indeed, since the inverse of the wealth is an anytime-valid p-value, it would yield a p-value of $1/(T+1)$, the best possible in this case.
@@ 2.1 Connection to Besag-Clifford
The aggressive strategy yields an anytime-valid generalization of the sequential permutation test by Besag and Clifford [3], when their method is stopped after one loss. Define the stopping time $\gamma(h,T)$ that stops after $h$ losses or after sampling $T$ (possibly infinite) permutations, whichever happens first. Recalling our notation $L_{t}$ for the number of losses after $t$ steps, Besag and Clifford defined a single p-value $Q^{\mathrm{BC}}_{\gamma(h,T)}$ (as opposed to a p-process) as
$Q^{\mathrm{BC}}_{\gamma(h,T)}=\begin{cases}h/\gamma(h,T),&L_{\gamma(h,T)}=h\\ (L_{\gamma(h,T)}+1)/(T+1),&\text{otherwise}.\end{cases}$ (8)
The p-value $Q^{\mathrm{BC}}_{\gamma(h,T)}$ is valid only at time $\gamma(h,T)$ and not at any other stopping time. While Besag and Clifford mentioned the possibility of $T=\infty$, they mainly focused on $T<\infty$ and most follow-up work ignored $T=\infty$ entirely. For this reason, we briefly discuss the case $T=\infty$ in Appendix C.
@@
Since our test yields a p-value at any stopping time, it makes sense to ask what the p-process (5) defined by the aggressive strategy $(Q^{\mathrm{agg}}_{t})_{t\in\mathbb{N}}$ equals at time $\gamma(1,T)$. The fascinating answer is that it equals the Besag-Clifford p-value at that stopping time:
$Q^{\mathrm{agg}}_{\gamma(1,T)}=Q^{\mathrm{BC}}_{\gamma(1,T)}.$
@@
The above equality has two key implications. First, it clearly leads to the interpretation that our aggressive betting strategy yields a p-process that is an anytime-valid generalization of the Besag-Clifford p-value that stops after one loss. (Our p-process recovers the latter p-value exactly at time $\gamma(1,T)$, but remains valid at all other stopping times.) Second, it is well known that $Q^{\mathrm{BC}}_{\gamma(1,T)}$ is in fact an exact p-value, meaning that even though it can only take on the values $1/1,1/2,1/3,\dots,1/(T+1)$, we have $\mathbb{P}(Q^{\mathrm{BC}}_{\gamma(1,T)}\leq c)=c$ for any $c$ in this support. Thus, the same property also holds for $Q^{\mathrm{agg}}_{\gamma(1,T)}$.
In Section 6, we provide an anytime-valid generalization of the Besag-Clifford p-value for general $h$, using a more sophisticated betting strategy.
2.2 Resampling risk
Since $Y_{1},Y_{2},\ldots$ are always exchangeable conditional on $Y_{0}$, it follows that $I_{1},I_{2},\ldots$ are also exchangeable. De Finetti’s representation theorem implies that $L_{t}/t\stackrel{{\scriptstyle a.s.}}{{\to}}Q_{\mathrm{lim}}$ for $t\to\infty$, where $Q_{\mathrm{lim}}$ is some random variable with values in $[0,1]$ and $I_{1},I_{2},\ldots$ are i.i.d. conditional on $Q_{\mathrm{lim}}$ with $\mathbb{P}(I_{1}=1|Q_{\mathrm{lim}})=Q_{\mathrm{lim}}$. If $H_{0}$ is true, then $Q_{\mathrm{lim}}$ follows a uniform distribution on $[0,1]$. If $H_{1}$ is true, then $Q_{\mathrm{lim}}$ is stochastically smaller than uniform. Consequently, the hypotheses $H_{0}$ and $H_{1}$ imply the following null $H_{0}^{\prime}$ and alternative $H_{1}^{\prime}$, respectively, formulated in terms of the indicators $I_{t}$ instead of the test statistics $Y_{t}$
$H_{0}^{\prime}:I_{1},I_{2},\ldots\text{ are i.i.d. }\mathrm{Ber}(q_{\mathrm{lim}})\text{ cond. on }Q_{\mathrm{lim}}=q_{\mathrm{lim}}\text{ with }Q_{\mathrm{lim}}\sim U[0,1],$
$H_{1}^{\prime}:I_{1},I_{2},\ldots\text{ are i.i.d. }\mathrm{Ber}(q_{\mathrm{lim}})\text{ cond. on }Q_{\mathrm{lim}}=q_{\mathrm{lim}}\text{ with }Q_{\mathrm{lim}}\text{ stochastically smaller than }U[0,1].$
The limiting permutation p-value $Q_{\mathrm{lim}}$ can also be interpreted as the p-value we would choose if we knew the distribution of $Y_{0}$ under $H_{0}$. Hence, a reasonable objective in the Monte Carlo test literature is to find a test $\phi\in\{0,1\}$ that bounds or minimizes the probability of obtaining a different decision than $Q_{\mathrm{lim}}$. This is called resampling risk [9], and is defined by
$\mathrm{RR}_{q_{\mathrm{lim}}}(\phi)=\begin{cases}\mathbb{P}(\phi=0|Q_{\mathrm{lim}}=q_{\mathrm{lim}}),&q_{\mathrm{lim}}\leq\alpha\\ \mathbb{P}(\phi=1|Q_{\mathrm{lim}}=q_{\mathrm{lim}}),&q_{\mathrm{lim}}>\alpha.\end{cases}$
Note that only the randomness induced by the resampling mechanism is relevant for the resampling risk $\mathrm{RR}_{q_{\mathrm{lim}}}$, since $Q_{\mathrm{lim}}=q_{\mathrm{lim}}$ is fixed. A small resampling risk guarantees that running the same permutation test (with different random seed) on the same data will likely yield the same result. Gandy [11] introduced an algorithm that uniformly bounds $\mathrm{RR}_{q_{\mathrm{lim}}}$ for all $q_{\mathrm{lim}}\in[0,1]$ by arbitrary small $\epsilon>0$ and which stops after a finite number of steps if $q_{\mathrm{lim}}\neq\alpha$. In this paper, we are mainly concerned with the case of $q_{\mathrm{lim}}\leq\alpha$, since our tests are valid by construction (which Gandy’s are not) such that additional rejections can be seen as power improvement and do not pose a problem for validity.
The aggressive strategy leads to a resampling risk of $\mathrm{RR}_{q_{\mathrm{lim}}}(\mathbbm{1}\{Q^{\mathrm{agg}}_{\tau}\leq\alpha\})=1-(1-q_{\mathrm{lim}})^{\lceil\frac{1}{\alpha}-1\rceil}\approx(q_{\mathrm{lim}}-\alpha q_{\mathrm{lim}})/\alpha$ for $q_{\mathrm{lim}}\leq\alpha$, which is greater than zero for all $q_{\mathrm{lim}}\neq 0$. Note that $q_{\mathrm{lim}}$ never equals exactly 00 (since there is a positive probability of drawing the identity) and therefore the resampling risk of the aggressive strategy is never zero. In Section 5, we introduce a “binomial mixture strategy” with zero resampling risk for all $q_{\mathrm{lim}}\in[0,c)$, where $c<\alpha$ can be chosen arbitrarily close to $\alpha$. Along the way, we introduce a “binomial strategy” with a zero resampling risk for all $q_{\mathrm{lim}}\in[0,\alpha/C)$ for an (explicit) constant $C$.
Note that the resampling risk of the permutation p-value and Besag-Clifford method equals zero only if $q_{\mathrm{lim}}\in\{0,1\}$ — and therefore is practically never zero. To the best of our knowledge, our methods are the first yielding zero resampling risk for a nontrivial set of $q_{\mathrm{lim}}$ values in the literature (whether using anytime-valid tests or not).
3 The oracle log-optimal strategy
Suppose we knew the true distribution of $I_{t}|I_{1}^{t-1}$, $t\in\mathbb{N}$, how should we place our bets $B_{t}$ to maximize the wealth $W_{T}$? This question is not easy to answer, since $W_{T}$ is a random variable and thus it is not clear in which sense it should be maximized. Intuitively, one could think of maximizing $\mathbb{E}[W_{T}]$. However, in this case we may end up with a strategy that leads to low or even zero wealth in many cases and to a very large wealth in a few cases. This seems not desirable. In order to avoid this, in similar problems, such as in portfolio theory or betting on horse races [6], the objective is to maximize the expected logarithmic wealth:
$\mathbb{E}[\log(W_{T})]=\sum_{t=1}^{T}\mathbb{E}[\log(B_{t}(I_{t}))].$ (9)
This is also called the Kelly criterion [16, 4]. The expected log-wealth has now become the de facto standard measure of performance in testing by betting; see [31, 13, 43, 26]. Despite not being the only option, we see no particular reason to deviate from it, so we adopt this perspective going forward in guiding our design of good betting strategies. However, it should be noted that log-optimality does not imply optimality of the corresponding p-process (5).
The log-wealth can be maximized by maximizing $\mathbb{E}[\log(B_{t}(I_{t}))|I_{1}^{t-1}]$ at each step $t\in\mathbb{N}$. Note that we need to condition on the past, since we are allowed to choose our betting strategy based on the previous loss indicators. A betting strategy that maximizes (9) is called a log-optimal strategy. One main reason to consider the logarithmic wealth is that the sum offers nice asymptotic behavior. For example, if the considered random variables $Z_{1},\ldots,Z_{T}$ are i.i.d., the strong law of large numbers implies that
$\frac{1}{T}\sum_{t=1}^{T}\log(Z_{1})\to\mathbb{E}[\log(Z_{1})]\text{ with probability }1.$
The loss indicators $I_{1},I_{2},\ldots$ are exchangeable under the null and the alternative and therefore the vector $I_{1}^{t-1}$ does not contain more information about $I_{t}$ than $L_{t-1}=\sum_{i=1}^{t-1}I_{i}$. Therefore, we replace $I_{1}^{t-1}$ by $L_{t-1}$ in the condition $\mathbb{P}(I_{t}=1|I_{1}^{t-1})$ from now on. In the following proposition we show a basic result that characterizes the log-optimal strategy.
Proposition 3.1.
Let $p_{t}^{*}=\mathbb{P}(I_{t}=1|L_{t-1})$, $t\in\mathbb{N}$. Then the (unique) log-optimal strategy is given by
$B_{t}^{*}(0)=(1-p_{t}^{*})\frac{t+1}{t-L_{t-1}}\quad\text{and}\quad B_{t}^{*}(1)=p_{t}^{*}\frac{t+1}{L_{t-1}+1.}$ (10)
@@ 3 The oracle log-optimal strategy
Let $b_{t}=B_{t}(1)(1+L_{t-1})/(t+1)$ for some betting strategy $B_{t}$. We obtain by Gibbs’ inequality
$\mathbb{E}[\log(B_{t}(I_{t}))|L_{t-1}]$
$=p_{t}^{*}\log(B_{t}(1))+(1-p_{t}^{*})\log(B_{t}(0))$
$=p_{t}^{*}\log\left(\frac{t+1}{1+L_{t-1}}\right)+(1-p_{t}^{*})\log\left(\frac{t+1}{t-L_{t-1}}\right)+p_{t}^{*}\log(b_{t})+(1-p_{t}^{*})\log(1-b_{t})$
$\leq p_{t}^{*}\log\left(\frac{t+1}{1+L_{t-1}}\right)+(1-p_{t}^{*})\log\left(\frac{t+1}{t-L_{t-1}}\right)+p_{t}^{*}\log(p_{t}^{*})+(1-p_{t}^{*})\log(1-p_{t}^{*})$
$=\mathbb{E}[\log(B_{t}^{*}(I_{t}))|L_{t-1}],$
with an equality iff $b_{t}=p_{t}^{*}$. ∎
Proposition 3.1 shows that the log-optimal strategy is to choose the bets proportional to the probabilities $p_{t}^{*}$ and $1-p_{t}^{*}$. This strategy is also called Kelly gambling [16, 4]. Using the limiting permutation p-value $Q_{\mathrm{lim}}$, we can also rewrite $p_{t}^{*}$ as
$p_{t}^{*}=\mathbb{P}(I_{t}=1|L_{t-1})=\mathbb{E}[\mathbb{P}(I_{t}=1|Q_{\mathrm{lim}})|L_{t-1}]=\mathbb{E}[Q_{\mathrm{lim}}|L_{t-1}],$
since $I_{1},I_{2},\dots$ are i.i.d. conditional on $Q_{\mathrm{lim}}$ with $\mathbb{P}(I_{t}=1|Q_{\mathrm{lim}})=Q_{\mathrm{lim}}$ (see Section 2.2). Hence, the log-optimal strategy tries to learn the realized value of $Q_{\mathrm{lim}}$ on our data set $X_{0}$ over time. If we would know the distribution of $Q_{\mathrm{lim}}$, then we could apply the log-optimal strategy. For example, $Q_{\mathrm{lim}}$ follows a uniform distribution on $[0,1]$ under $H_{0}$. Hence, $Q_{\mathrm{lim}}|L_{t-1}=\ell$ follows a Beta distribution with parameters $\alpha=\ell+1$ and $\beta=t-\ell$ under $H_{0}$, which implies that $p_{t}^{*}=\mathbb{E}_{H_{0}}[Q_{\mathrm{lim}}|L_{t-1}]=(L_{t-1}+1)/(t+1)$. Therefore, the log-optimal strategy under $H_{0}$ is the powerless passive betting strategy introduced in Section 2. This makes sense, as we should not be able gain wealth under the null. Under the alternative, $Q_{\mathrm{lim}}$ is stochastically smaller than uniform. Therefore, $B_{t}^{*}(0)\geq B_{t}^{*}(1)$, meaning our wealth increases in case of a win (and decreases otherwise). Without additional information about the distribution of $Q_{\mathrm{lim}}$ under $H_{1}$, it is not possible to pin down the log-optimal bet further.
In Appendix D, we characterize the asymptotic behavior of the log-optimal strategy which turns out to be quite different from the standard behavior observed in betting games. In particular, we prove that $B_{t}^{*}(I_{t})\to 1$ almost surely for $t\to\infty$, implying that we can only accumulate wealth very slowly at late stages of the testing process.
4 The binomial strategy
We now use the oracle results from Section 3 to define powerful (fully data-driven) strategies. Therefore, we choose $B_{t}(r)$ as the log-optimal strategy from (10), where we replace
$p_{t}^{*}=\mathbb{P}(I_{t}=1|L_{t-1})=\mathbb{E}[Q_{\mathrm{lim}}|L_{t-1}]$
by some reasonable hyperparameter $p_{t}$. Recall from the previous section that if we would know the distribution of $Q_{\mathrm{lim}}$, the log-optimal bet would be to set $p_{t}=\mathbb{E}[Q_{\mathrm{lim}}|L_{t-1}]$. However, such information is usually not available as permutation tests are used in non-parametric settings and the alternative $H_{1}$ is composite. An intuitive practical approach is therefore to choose some reasonable “working prior” density $f$ for $Q_{\mathrm{lim}}$ instead and calculate the log-optimal betting strategy with respect to $f$, meaning to set $p_{t}=\mathbb{E}_{f}[Q_{\mathrm{lim}}|L_{t-1}]$. We call this the mimicked log-optimal strategy. Note that if $f$ is not the true density of $Q_{\mathrm{lim}}$, then the mimicked log-optimal strategy is still valid (meaning that it results in valid e-values and p-values at any stopping time), since it is valid for all $p_{t}\in[0,1]$ due to (2). However, it would not be log-optimal in this case.
The mimicked log-optimal strategy requires to calculate the posterior distribution $f_{Q_{\mathrm{lim}}|L_{t-1}}$ of $Q_{\mathrm{lim}}$ given the working prior $f$ and the previous number of losses $L_{t-1}$ to then set $p_{t}=\mathbb{E}_{f_{Q_{\mathrm{lim}}}|L_{t-1}}[Q_{\mathrm{lim}}]$ at each step. This is computationally demanding and thus not well suited for our sequential permutation tests. For this reason, we begin with considering a constant working prior $Q_{\mathrm{lim}}=p$ for some fixed $p\in[0,1]$ in this section, which does not require any updating. Note that this working prior never describes the true distribution of $Q_{\mathrm{lim}}$, as it assumes that the observed test statistic $Y_{0}$ is fixed. However, it is a simple starting point and already bears interesting results. In particular, it yields a simple closed form for the wealth. In Section 5, we show how this can be used to derive closed forms for more complex working priors $f$, thus avoiding the computational burden of updating the posterior distribution. For interpretation, one can think of the working prior $Q_{\mathrm{lim}}=p$ as our choice if we cheated and peeked at the actually realized value $y_{0}$ of $Y_{0}$ before choosing a betting strategy. In this case the only randomness comes from the random sampling mechanism.
4.1 Closed-form wealth of the binomial strategy
The binomial strategy is defined by (10) with some user-chosen constant $p\in[0,1]$ plugged in place of the unknown $p_{t}^{*}$. The wealth of the binomial strategy after $T$ permutations has a simple closed form, as we will show in the following.